The January 1, 2011 statement of financial position of Skittle Company at book and market values is as follows:

Book Value

Fair Value

Current Assets

P    800,000

P   750,000

Property and Equipment (net)

900,000

1, 000,000

Total Assets

P 1,700,000

P1,750,000

Current Liabilities

P    300,000

P   300,000

Long-term Liabilities

500,000

460,000

Ordinary Share Capital, P1 par

100,000

Share Premium

200,000

Accumulated Profits

600,000

Total Liabilities and Shareholders’ equity

P1,700,000

Polypeptide Company paid P950,000 in cash for 90% of Skittle Company’s ordinary share capital.

Required:

  1. Prepare a journal entry on Polypeptide’s books to record the acquisition of the Skittles share.
  2. Prepare a schedule of the allocation of the difference between the consideration given and the book value of interest acquired from Skittle Company.
  3. Prepare the elimination entries for the consolidated statements worksheet.
Answer #1
Amounts in P
a Acquisition Price paid =                      950,000
% Holding in Skittle 90%
b Implied Value of Skittle Net Assets =950000/90%=                   1,055,556
c Non Controlling Interest at Fair Value =                      105,556
Schedule of Allocation of Difference between Consideration Given and Book Value
Net Assets Acquired Book Value Fair Value Difference
Current Assets                      800,000                 750,000                (50,000)
Net P& E                      900,000              1,000,000                100,000
d Total Assets                   1,700,000              1,750,000                  50,000
Current Liab                      300,000                 300,000                          -  
LT Liab                      500,000                 460,000                (40,000)
e Total Liabilities                      800,000                 760,000                (40,000)
f Net Assets d-e=                      900,000                 990,000                  90,000
Ans b.
Allocation of Difference 90% Parent 10% NCI Total
m Purchase Consideration and Total Implied value                      950,000                 105,556            1,055,556
Less Book Value of Equity Acquired
Ordinary Share Capital P1 par                        90,000                   10,000                100,000
Share Premium                      180,000                   20,000                200,000
Accumulated Profit                      540,000                   60,000                600,000
n Total Book Value                      810,000                   90,000                900,000
o Difference between Impied and Book Value=m-n                      140,000                   15,556                155,556
Allocation of Difference
Current Assets                      (45,000)                   (5,000)                (50,000)
Net P&E                        90,000                   10,000                100,000
LT Liability                        36,000                     4,000                  40,000
Good Will Recognized                        59,000                     6,556                  65,556
Total Difference between Impied and Book Value=m-n                      140,000                   15,556                155,556
Ans a.
Journal Entry for Acquisition of Skittle in the Book of Polypeptide
Account Title Dr $ Cr $
Current Assets                      750,000
Net P& E                   1,000,000
Goodwill                        65,556
Current Liabilities                 300,000
LT Liabilities                 460,000
NCI (at FV of Net Assets)                 105,556
Cash                 950,000
                  1,815,556              1,815,556
Ans c. Elimination Entry
Account Title Dr $ Cr $
Ordinary Share Capital P1 par( Skittle)                      100,000
Share Premium(Skittle)                      200,000
Accumulated Profit(Skittle)                      600,000
Investment                 950,000
NCI                 105,556
Difference between Impied and Book Value                      155,556